Use appropriate software to sketch each of the following surfaces and identify it.

4(x+1)2+2(y-2)2+(z-1)2=1


1
Expert's answer
2020-12-28T14:54:49-0500

"4(x+1)^2+2(y-2)^2+(z-1)^2=1"

"\Rightarrow
\frac{(x+1)^2}{1/4}+\frac{(y-2)^2}{1/2}+\frac{(z-1)^2}{1}=1"


The above equation represents an Ellipsoid.

Graph :-

x-y Trace :-



x-z Trace :-

y-z Trace :-